# SkyScenes: A Synthetic Dataset for Aerial Scene Understanding

### v1

@@ B.2 Class Descriptions
ground: any horizontal ground-level structures that does not match any other category
@@
B.3 Depth & Instance Segmentation
We also provide depth and instance segmentation annotations along with semantic segmentation annotations. Both the depth and instance segmentation sensors are mounted alongside the RGB camera and semantic segmentation sensors. Depth is stored in the format which provides better results for closer objects. We also provide depth-aided semantic segmentation results (training details in Sec. C.3).
Appendix C Experiment Details
C.1 Class Merging Details
| Syn UAVid | UAVid | SkyScenes | SynDrone |
| 1 clutter | clutter | unlabelled | unlabelled |
| other | other | ||
| traffic sign | traffic sign | ||
| rail track | rail track | ||
| guard rail | guard rail | ||
| traffic light | traffic light | ||
| static | static | ||
| dynamic | dynamic | ||
| ground | ground | ||
| sidewalk | sidewalk | ||
| fence | fence | ||
| sky | sky | ||
| bridge | bridge | ||
| water | water | ||
| pole | pole | ||
| 2 building | building | building | building |
| wall | wall | ||
| 3 road | road | road | road |
| roadline | roadline | ||
| 4 vegetation | vegetation | vegetation | vegetation |
| 5 low vegetation | terrain | terrain | terrain |
| 6 person | person | pedestrian | pedestrian |
| rider | rider | ||
| motorcycle | motorcycle | ||
| bicycle | bicycle | ||
| 7 vehicle | car | car | car |
| truck | truck | ||
| bus | bus | ||
@@ C.1 Class Merging Details
| Syn ICG Drone | ICG Drone | SkyScenes | SynDrone |
| 1 other | obstacle | unlabelled | unlabelled |
| dog | other | other | |
| conflicting | traffic sign | traffic sign | |
| ar-marker | sky | sky | |
| unlabelled | bridge | bridge | |
| gurad rail | guard rail | ||
| traffic light | traffic light | ||
| static | static | ||
| dynamic | dynamic | ||
| 2 fence | fence | fence | fence |
| 3 pole | fence pole | pole | pole |
| 4 vegetation | tree | vegetation | vegetation |
| bald-tree | |||
| 5 building | wall | wall | wall |
| roof | building | building | |
| door | |||
| window | |||
| 6 water | water | water | water |
| pool | |||
| 7 bicycle | bicycle | bicycle | bicycle |
| motorcycle | motorcycle | ||
| 8 vehicle | car | car | car |
| truck | truck | ||
| bus | bus | ||
| train | |||
| 9 person | person | pedestrian | pedestrian |
| rider | rider | ||
| 10 paved area | paved area | ground | ground |
| sidewalk | sidewalk | ||
| road | road | ||
| roadline | roadline | ||
| rail track | rail track | ||
| 11 terrain | rocks | terrain | terrain |
| gravel | |||
| dirt | |||
| vegetation | |||
@@
SyntheticReal. As noted in Sec of the main paper, since different real-world datasets have different class vocabularies and definitions, for our SyntheticReal semantic segmentation experiments, we adapt the class-vocabulary of the synthetic source dataset (SkyScenes, SynDrone) to that of the target real dataset (UAVid, AeroScapes, ICG Drone). This is done using a class-merging scheme based on the class-vocabularies and after visually inspecting dataset annotations. We provide the class-merging schemes used for both SkyScenes and SynDrone [25] (synthetic) across real counterparts UAVid [19] (in Table. 8), AeroScapes [22] (in Table. 9) and ICG Drone [14] (in Table. 10).
SkyScenes Diagnostic Experiments. To assess the sensitivity of trained models to different factors – weather, time of day, height, pitch, etc.– we train models on different SkyScenes variations and evaluate them on held-out-variations. For these experiments, we recude the SkyScenes vocabulary to a reasonable subset of classes (consistent with the widely used Cityscapes [5] palette) – road, sidewalk, building, wall, fence, pole, traffic light, traffic sign, vegetation, water, sky, pedestrian, rider, cars, truck, bus, roadline, motorcycle, bicycle and an ignore class.
C.2 Training, Validation and Test Splits
C.2.1 Synthetic Real Experiments
@@
% ( images) each for validation and testing.
SynDrone. SynDrone has images per town (across Carla towns) for each of the
combinations,
resulting in a total of images per
combination.
We use of these datapoints for training models, with kept aside for testing. The datapoints selected for training and testing are kept consistent with the one reported in SynDrone [25].
C.2.2 SkyScenes Diagnostic Experiments
Weather & Daytime Variations. For each weather variation we sample evenly across the towns and
combinations (excluding ),
resulting in a total of images.
We use
% ( images) of these
for training models, and remaining % ( images) each for validation and testing. We evaluate the model on the same weather variation it was trained on and select the model with the best mIoU score for further evaluations on other weather variations.
@@ C.2.2 SkyScenes Diagnostic Experiments
C.3 Training Details
Semantic Segmentation. For our semantic segmentation experiments we use both CNN – DeepLabv2 [3] (ResNet-101 [10] backbone) – and vision transformer – DAFormer [11] (with HRDA [12] source training; MiT-B5 [36] backbone) – based semantic segmentation architectures. Note that we utilize only the DAFormer architecture to perform source-only training for our experiments. Following [13], we enable rare class sampling [11] and use Imagenet feature-distance for our thing classes during training. All models are trained using the AdamW [18] optimizer coupled with a polynomial learning rate scheduler with an initial learning rate of . Each model is trained for k iterations with a batch size of . For our fine-tuning experiments, we use an initial learning rate of .
@@ C.4 Evaluation Details
Appendix D Results
D.1 SyntheticReal Aligned Data Selection
@@
| other | fence | pole | vegetation | building | water | bicycle | vehicle | person | paved area | terrain | Avg | Avg | ||
@@
| Target | |||||||||
| 21 Target | |||||||||
| 22 Target | |||||||||
| 23 Target | |||||||||
| 24 Target | |||||||||
| 25 Target | |||||||||
| Training | Data Size | Synthetic mIoU () | |||||||
| SkyScenes | |||||||||
| 1 FT | |||||||||
| 2 FT | |||||||||
| 3 FT | |||||||||
| 4 FT | |||||||||
| 5 FT | |||||||||
| 6 JT | 62.97 | ||||||||
| 7 JT | 67.58 | ||||||||
| 8 JT | 70.20 | ||||||||
| 9 JT | 71.83 | ||||||||
| 10 JT | |||||||||
| SynDrone | |||||||||
| 11 FT | |||||||||
| 12 FT | |||||||||
| 13 FT | |||||||||
| 14 FT | |||||||||
| 15 FT | |||||||||
| 16 JT | |||||||||
| 17 JT | |||||||||
| 18 JT | |||||||||
| 19 JT | |||||||||
| 20 JT | |||||||||
@@
D.2 SkyScenes Real Data Experiments
In addition to zero-shot transfer to real data, we also show how SkyScenes is useful as additional training data when labeled real-world data is available. In Fig. 8 and Fig. 9, we compare the performance of DeepLabv2 [3] for SkyScenes UAVid [19] and for SkyScenes AeroScapes [22] trained only using of UAVid [19] and AeroScapes [22] training images respectively with counterparts that were either pretrained using SkyScenes data or additionally supplemented with SkyScenes data at training time. In Fig. 10 and Fig. 11 we make a similar comparison with the DAFormer [11] architecture. In low-shot regimes (when little “real” world data is available), SkyScenes data (either explicitly via joint training or implicitly via finetuning) is beneficial in improving recognition performance. We find this to be especially beneficial for under-represented classes in aerial imagery (such as humans and vehicles). In Table. 13(a) and Table. 13(c) we present similar fine-grained (per-class) comparison of SkyScenes with SynDrone for a DeepLabv2 model when real data is available for training – via Target-Only, Finetuning or Joint-Training. Tables. 13(b) and 13(d) show similar comparisons using the DAFormer [11] model.
D.3 SkyScenes Diagnostic Experiments
Similar to Table 6 (d) in the main paper, in Table 15, we assess broader sensitivity of models by training DAFormer [11] models across all (total ) settings and evaluate them across the same conditions in SkyScenes. Table 15 (a) models trained on are representative of one extreme of the range (both lowest height and pitch values) – we notice that this model is extremely sensitive to height variations. From Tables 15 (a), (e) and (i), we can deduce that due to the high variability in perspective between and other conditions, models trained on do not generalize well to other values. On the other hand, Tables. 15 (b), (c), (d), (f), (g) and (h) models trained on oblique perspectives are better at generalizing to other pitch conditions.
@@ D.3 SkyScenes Diagnostic Experiments
D.4 Qualitative examples
In Fig. 12, 13 and 14, we show qualitative examples of predictions made by SkyScenes and SynDrone trained DeepLabv2 and DAFormer models on UAVid, AeroScapes, and ICG Drone respectively. In Fig. 15 and 16, we show how predictions are impacted by changing SkyScenes conditions.

### v2

@@ 0.B.2 Class Descriptions
ground: any horizontal ground-level structures that do not match any other category
@@
0.B.3 Depth & Instance Segmentation
We also provide depth and instance segmentation annotations along with semantic segmentation annotations. Both the depth and instance segmentation sensors are mounted alongside the RGB camera and semantic segmentation sensors. Depth is stored in the format which provides better results for closer objects. We also provide depth-aided semantic segmentation results (training details in Sec. 0.C.4).
0.B.4 SkyScenes vs Real Characteristics
| FID Score () | |||
| UAVid | AeroScapes | ICG Drone | |
| 1 SynDrone | |||
| 2 SkyScenes | 8.78 | 7.66 | 5.77 |
In Fig. 7, we compare perceptual similarity (in terms of FID [15]) of synthetic images from SkyScenes and SynDrone [42] with real datasets (UAVid [27], AeroScapes [34], ICG Drone [20]). We find SkyScenes images are closer to real data distributions than SynDrone [42]. For semantic comparisons, we show class-distribution comparisons (per-class pixel frequencies) between SkyScenes, SynDrone [42] and UAVid [27] in Fig 7. SkyScenes images display higher distribution of tail classes(vehicles, humans) compared to SynDrone [42], and shows better alignment with the real data distribution (UAVid [27]).
Appendix 0.C Experiment Details
0.C.1 Class Merging Details
| Syn UAVid | UAVid | SkyScenes | SynDrone | VALID | SynthAer |
| 1 clutter | clutter | unlabelled | unlabelled | background | sky |
| sidewalk | sidewalk | pavement | footpath | ||
| fence | fence | fence | |||
| bridge | bridge | bridge | |||
| water | water | water | |||
| traffic light | traffic light | traffic light | |||
| other | other | sign | |||
| traffic sign | traffic sign | land | |||
| rail track | rail track | tunnel | |||
| guard rail | guard rail | pool | |||
| static | static | stones | |||
| dynamic | dynamic | pierruble | |||
| ground | ground | chair | |||
| sky | sky | ice | |||
| pole | pole | ship | |||
| plane | |||||
| harbor | |||||
| lamp | |||||
| bus stop | |||||
| powerline | |||||
| garbage bin | |||||
| other low obstacle | |||||
| other high obstacle | |||||
| 2 building | building | building | building | building | building |
| wall | wall | wall | |||
| 3 road | road | road | road | road | road |
| roadline | roadline | ||||
| 4 vegetation | vegetation | vegetation | vegetation | tree | tree |
| 5 low vegetation | terrain | terrain | terrain | other plant | vegetation |
| 6 person | person | pedestrian | pedestrian | person | |
| rider | rider | ||||
| motorcycle | motorcycle | ||||
| bicycle | bicycle | ||||
| 7 vehicle | car | car | car | small vehicle | car |
| truck | truck | large vehicle | |||
| bus | bus | ||||
@@ 0.C.1 Class Merging Details
| Common Scheme | UAVid | AeroScapes | ICG Drone | SkyScenes |
| 1 clutter | clutter | background | unlabeled | unlabeled |
| sky | dirt | sky | ||
| grass | pole | |||
| gravel | sidewalk | |||
| water | traffic sign | |||
| pool | other | |||
| rock | ground | |||
| fence-pole | guard rail | |||
| dog | traffic light | |||
| ar-marker | static | |||
| obstacle | dynamic | |||
| conflicting | water | |||
| terrain | ||||
| 2 road | road | road | paved area | road |
| roadline | ||||
| rail track | ||||
| 3 nature | vegetation | vegetation | vegetation | vegetation |
| tree | tree | |||
| bald tree | ||||
| 4 person | human | person | person | pedestrian |
| bicycle | bicycle | bicycle | ||
| rider | ||||
| motorcycle | ||||
| 5 vehicle | car | vehicle | car | car |
| truck | ||||
| bus | ||||
| 5 construction | building | building | roof | building |
| wall | wall | |||
| fence | fence | |||
| window | bridge | |||
| door |
@@
| Syn ICG Drone | ICG Drone | SkyScenes | SynDrone | VALID | SynthAer |
| 1 other | obstacle | unlabelled | unlabelled | background | sky |
| dog | other | other | animal | ||
| conflicting | traffic sign | traffic sign | sign | ||
| ar-marker | sky | sky | ice | ||
| unlabelled | bridge | bridge | bridge | ||
| gurad rail | guard rail | stones | |||
| traffic light | traffic light | traffic light | |||
| static | static | pierruble | |||
| dynamic | dynamic | garbage bin | |||
| plane | |||||
| harbor | |||||
| land | |||||
| chair | |||||
| ship | |||||
| lamp | |||||
| tunnel | |||||
| bus stop | |||||
| powerline | |||||
| garbage bin | |||||
| other low obstacle | |||||
| other high obstacle | |||||
| 2 fence | fence | fence | fence | fence | |
| 3 pole | fence pole | pole | pole | ||
| 4 vegetation | tree | vegetation | vegetation | tree | tree |
| bald-tree | other plant | vegetation | |||
| 5 building | wall | wall | wall | wall | |
| roof | building | building | building | ||
| door | |||||
| window | |||||
| 6 water | water | water | water | water | |
| pool | pool | ||||
| 7 bicycle | bicycle | bicycle | bicycle | ||
| motorcycle | motorcycle | ||||
| 8 vehicle | car | car | car | small vehicle | vehicle |
| truck | truck | large vehicle | |||
| bus | bus | ||||
| train | |||||
| 9 person | person | pedestrian | pedestrian | person | |
| rider | rider | ||||
| 10 paved area | paved area | road | road | road | road |
| sidewalk | sidewalk | pavement | footpath | ||
| ground | ground | ||||
| roadline | roadline | ||||
| rail track | rail track | ||||
| 11 terrain | rocks | terrain | terrain | land | |
| gravel | |||||
| dirt | |||||
| vegetation | |||||
@@
SyntheticReal. As noted in Sec. of the main paper, since different real-world datasets have different class vocabularies and definitions, for our Synthetic
Real semantic segmentation experiments, we adapt the class-vocabulary of the synthetic source dataset (SkyScenes, SynDrone, VALID, SynthAer) to that of the target real dataset (UAVid, AeroScapes, ICG Drone). This is done using a class-merging scheme based on the class-vocabularies and after visually inspecting dataset annotations. We provide the class-merging schemes used for the synthetic datasets (SkyScenes, SynDrone [42], VALID [5] and SynthAer [47]) across real counterparts UAVid [27] (in Table. 7), AeroScapes [34] (in Table. 9) and ICG Drone [20] (in Table. 10). We also include a coarser common merging scheme for SkyScenes and all the real datasets UAVid, AeroScapes, and ICG Drone in Table. 8.
SkyScenes Diagnostic Experiments. To assess the sensitivity of trained models to different factors – weather, time of day, height, pitch, etc.– we train models on different SkyScenes variations and evaluate them on held-out-variati-ons. For these experiments, we reduce the SkyScenes vocabulary to a reasonable subset of classes (consistent with the widely used Cityscapes [8] palette) – road, sidewalk, building, wall, fence, pole, traffic light, traffic sign, vegetation, water, sky, pedestrian, rider, cars, truck, bus, roadline, motorcycle, bicycle and an ignore class.
0.C.2 Training, Validation and Test Splits
@@
% ( images) each for validation and testing. While creating train, val and test splits, we collect equal number of samples from each town by dividing each town-specific traversal sequence into segments: the initial % for training, the next % for testing, and the final % of the segment for validation. Moreover, within each split, we ensure that every viewpoint is accompanied by its different variations across weather, daytime, height, and pitch settings. This safeguards against any potential cross-contamination across different splits while ensuring fair representation and equal distributions of all variations.
SynDrone. SynDrone has images per town (across Carla towns) for each of the combinations,
resulting in a total of images per combination. We use of these data points for training models, with kept aside for testing. The data points selected for training and testing are kept consistent with the one reported in SynDrone [42].
VALID.VALID has a total of images spread across different height variations and different layout and daytime conditions all from the nadir perspective. For the combination, and , the total images are , and respectively which are split into for training and each for validation and testing. The data points selected for training and testing are kept consistent with the one reported in VALID [5]
SynthAer. The split of the dataset and the data points selected for training and testing are consistent with the one reported in SynthAer [47] with images for training, for validation and for testing.
0.C.3 SkyScenes Diagnostic Experiments
Weather & Daytime Variations. For each weather variation we sample evenly across the towns and combinations (excluding ),
resulting in a total of images. We use % ( images) of these for training models, and remaining % ( images) each for validation and testing. We evaluate the model on the same weather variation it was trained on and select the model with the best mIoU score for further evaluations on other weather variations.
@@ 0.C.3 SkyScenes Diagnostic Experiments
0.C.4 Training Details
Semantic Segmentation. For our semantic segmentation experiments we use both CNN – DeepLabv2 [4] (ResNet-101 [14] backbone) – and vision transformer – 1. DAFormer [16] (with HRDA [17] source training; MiT-B5 [55] backbone) and 2. Rein DINOv2 [54] – based semantic segmentation architectures. Note that we utilize only the DAFormer architecture to perform source-only training for our experiments. Following [18], we enable rare class sampling [16] and use Imagenet feature-distance for our thing classes during training. DeepLabv2 [4] and DAFormer [16] are trained using the AdamW [26] optimizer coupled with a polynomial learning rate scheduler with an initial learning rate of . For our fine-tuning experiments, we use an initial learning rate of . For Rein [54], we use initial learning rate of and for fine-tuning we use . Each model is trained for k iterations with a batch size of .
@@ 0.C.5 Evaluation Details
Appendix 0.D Results
0.D.1 Ground vs Aerial Performance Gap
Recent advancements in synthetic-to-real (syn-to-real) generalization have significantly closed the performance gap in ground imagery to about 20%, unlike in aerial imagery where the gap remains as high as 50%. In our study, we evaluated Rein DINOv2 [54] across ground-view settings using GTAVCityscapes and aerial-view settings using SkyScenes ICG. For ground views, the gap between GTAVCityscapes and Cityscapes [8] mIoU was 16.29%, demonstrating a mIoU of compared to . Aerial views showed a larger gap, with SkyScenes ICG mIoU at versus ICG mIoU at , resulting in a 50.53% gap. This highlights the need for focused efforts to address the larger performance discrepancies observed in aerial imagery.
@@ 0.D.1 Ground vs Aerial Performance Gap
| other | fence | pole | vegetation | building | water | bicycle | vehicle | person | paved area | terrain | Avg | Avg∗ | ||
@@
0.D.2 Advantage of SkyScenes over Real Datasets
Curating densely annotated real data under diverse conditions in a controlled / uncontrolled manner is prohibitively expensive. As a synthetic alternative, SkyScenes is well-suited for evaluating model sensitivity to changing conditions. Further, due to the lack of diverse annotated data, we observe that real aerial datasets are somewhat homogenous. Consequently, models trained on real data under specific conditions (unlike SkyScenes) struggle to generalize to differing conditions – Rein DINOv2 [54] models trained on SkyScenes generalize better to ICG Drone [20] ( mIoU) compared to ones trained on AeroScapes [34] ( mIoU) and UAVid [27] ( mIoU)666A common merging scheme was used as detailed in Table. 8
0.D.3 SyntheticReal Aligned Data Selection
0.D.4 SyntheticReal Additional Experiments
| (Source) | (Target) Real-World mIoU () | ||
| Synthetic | UAVid | AeroScapes | ICG Drone |
| 1 SynthAer | |||
| 2 VALID | |||
| 3 SynDrone | 54.92 | ||
| 4 SkyScenes | 43.96 | 28.10 | |
A proper comparison with other synthetic datasets on a recognition task would require similar classes, (which a subset of synthetic alternatives, such as MidAir [12], do not satisfy). Considering other desired attributes (outlined in Table.1 in the paper), we chose SynDrone [42] as the most aligned alternative for our experiments. For comprehensive analysis and additional comparisons with other synthetic alternatives across various SyntheticReal settings, we included comparisons with VALID [5], which features only nadir perspective, and SynthAer [47], which lacks human classes, in Table 14. From the results in Table 14 , it is evident that models trained on SkyScenes generalize better to real datasets compared to other synthetic alternatives.
0.D.5 SkyScenes Real Data Experiments
@@
| Training | Data Size | Synthetic mIoU () | |||||||
| SkyScenes | |||||||||
| 1 FT | |||||||||
| 2 FT | |||||||||
| 3 FT | |||||||||
| 4 FT | |||||||||
| 5 FT | |||||||||
| 6 JT | 62.97 | ||||||||
| 7 JT | 67.58 | ||||||||
| 8 JT | 70.20 | ||||||||
| 9 JT | 71.83 | ||||||||
| 10 JT | |||||||||
| SynDrone | |||||||||
| 11 FT | |||||||||
| 12 FT | |||||||||
| 13 FT | |||||||||
| 14 FT | |||||||||
| 15 FT | |||||||||
| 16 JT | |||||||||
| 17 JT | |||||||||
| 18 JT | |||||||||
| 19 JT | |||||||||
| 20 JT | |||||||||
| Target | |||||||||
| 21 Target | |||||||||
| 22 Target | |||||||||
| 23 Target | |||||||||
| 24 Target | |||||||||
| 25 Target | |||||||||
@@
| Training | Data Size | Synthetic mIoU () | |||||||
| SkyScenes | |||||||||
| 1 FT | 70.7 | ||||||||
| 2 FT | 73.99 | ||||||||
| 3 FT | |||||||||
| 4 FT | 76.44 | ||||||||
| 5 FT | 76.89 | ||||||||
| 6 JT | |||||||||
| 7 JT | |||||||||
| 8 JT | 75.08 | ||||||||
| 9 JT | |||||||||
| 10 JT | |||||||||
| Target | |||||||||
| 21 Target | |||||||||
| 22 Target | |||||||||
| 23 Target | |||||||||
| 24 Target | |||||||||
| 25 Target | |||||||||
In addition to zero-shot transfer to real data, we also show how SkyScenes is useful as additional training data when labeled real-world data is available. In Fig. 8 and Fig. 9, we compare the performance of DeepLabv2 [4] for SkyScenes UAVid [27] and for SkyScenes AeroScapes [34] trained only using of UAVid [27] and AeroScapes [34] training images respectively with counterparts that were either pretrained using SkyScenes data or additionally supplemented with SkyScenes data at training time. In Fig. 10 and Fig. 11 we make a similar comparison with the DAFormer [16] architecture, and in Fig. 12 with the Rein DINOv2 [54] architecture. In low-shot regimes (when little “real” world data is available), SkyScenes data (either explicitly via joint training or implicitly via finetuning) is beneficial in improving recognition performance. We find this to be especially beneficial for under-represented classes in aerial imagery (such as humans and vehicles). In Table. 15 and Table. 16 we present similar fine-grained (per-class) comparison of SkyScenes with SynDrone for a DeepLabv2 model when real data is available for training – via Target-Only, Finetuning or Joint-Training.Tables. 17 and 18 show similar comparisons using the DAFormer [16] model. Table. 19 shows similar comparison using the Rein DINOv2 [54] model.
0.D.6 SkyScenes Diagnostic Experiments
@@
Similar to Table 6 (d) in the main paper, in Table 20, we assess broader sensitivity of models by training DAFormer [16] models across all (total ) settings and evaluate them across the same conditions in SkyScenes.Table 20 (a) models trained on are representative of one extreme of the range (both lowest height and pitch values) – we notice that this model is extremely sensitive to height variations.Tables 20 (a), (e) and (i), we can deduce that due to the high variability in perspective between and other conditions, models trained on do not generalize well to other values. On the other hand, Tables. 20 (b), (c), (d), (f), (g) and (h) models trained on oblique perspectives are better at generalizing to other pitch conditions.
| Daytime | Test mIoU () | ||||
| Conditions | Vanilla | Motion Blur | Spatter | ||
| sev. = 3 | sev. = 5 | sev. = 3 | sev. = 5 | ||
| Noon | 83.54 | 74.31 (9.23) | 67.49 (16.05) | 78.42 (5.11) | 63.03 (20.51) |
| Sunset | 84.32 | 74.62 (9.70) | 68.08 (16.24) | 78.10 (6.22) | 68.62 (15.69) |
0.D.7 Visual Artifacts
Real-world images can have sensory artifacts beyond weather/daytime variations (UAV123 [30], VIRAT [35], UG2 [52]). Prior work (RobustNav [2]) shows it’s possible to simulate such corruptions with varying severities. Fig.0.D.6 provides examples of motion blur and spatter, highlighting how models trained on clean data degrade under these conditions (see Table. 0.D.6).
0.D.8 Auto labeling for Real Datasets
Fig.13 shows a sample semantic segmentation prediction of a random UAVid image using Grounded-SAM [41]. We find that most of the roads are absent from the prediction, no humans are predicted, and only a few instances of vehicles, buildings, and trees are identified correctly. While auto-labeling using foundation models has potential, we believe there is still room for improvement in specific applications (such as aerial).
0.D.9 Qualitative examples
Appendix 0.E Limitations
Depth of distant objects in SkyScenes is imperfect, akin to real-world conditions. As evident from Fig. 5 in the paper, depth observations struggle to discern finer details of high-altitude viewpoints. We hypothesize that improvement from depth is helpful only for objects close to the camera or lower altitude viewpoints.
Appendix 0.F Future Work
We plan on updating SkyScenes with evolving considerations for real-world aerial scene-understanding – improved realism, additional anticipated edge cases – as more and more features are supported in the underlying simulator and provide additional support for 3D perception tasks (see Fig. 19.)